Physical Chemistry

Rate of formation of SO3 according to the reaction 2SO2 + O2 – 2SO2 is 1.6*10^-3 kg/min.Hence rate at witch SO2 reacts is-
a.1.28 * 10^-3 b.1.6 * 10^-3

Tanusri Bose
7 Years agoGrade

1 Answer

Arun
7 Years ago
rate of reaction= 1/2* Rate of disappearance of SO2 =1/2*Rate of formation of SO3
​=1/2* Rate of disappearance of SO2 = 1/2*1.6×10^-3
therefore, Rate of disappearance of SO2= 
2 × 8
Rate of disappearance of SO2= 1.6 *10^-3kg/min